For y ∈ Z, prove that 9y² + 3y − 2 is even.


1
Expert's answer
2022-04-26T00:19:27-0400

Consider two cases:

  1. Suppose that "y" is odd. It means, that we can present "y" as: "y=2n+1", where "n\in\mathbb{Z}". We get: "9(2n+1)^2+3(2n+1)-2=9(4n^2+4n+1)+6n+1=36n^2+42n+10". The latter number is even
  2. Suppose that "y" is even. It means that we can present "y" as: "y=2n". We receive: "36n^2+6n-2." The latter number is even.

Thus, we have shown that the number "9y^2+3y-2" is even.